# Lucas Resende CREST, ENSAE, Institut Polytechnique de Paris, France Guillaume Lecué ESSEC Business School, France Lionel Wilner CREST, ENSAE, Institut Polytechnique de Paris, France Philippe Choné CREST, ENSAE, Institut Polytechnique de Paris, France

### v1

@@ B.2.1 Hájek projection of the gradient of the loss function at $\beta_{\star}$
The main difficulty in proving a CLT for $U_{N}$ is that it is not a sum of independent variables. However, we prove in this section, as in graham2017econometric; jochmans2018semiparametric, that it can be well approximated (in a sense given in Proposition 2 below) by its Hájek projection - which is a sum of independent variables.
@@
is achieved by $U_{N}^{*}$. Note that the $(\sigma(X,Y_{\bf i}))_{{\bf i}}$ are not independent in general, however, they are independent conditionally on $X$ because all edges weights are independent conditionally on $X$ in model assumption (1). That is the reason why we are considering the Hájeck projection conditionally to $X$. However, because of this conditioning, we cannot directly apply Lemma 11.10 from MR1652247 but we can adapt its proof to our context. We therefore follow the proof strategy of Lemma 11.10 from MR1652247 to show that $U_{N}^{*}$ is solution to (B.5).
@@ B.2.2 Asymptotic normality of the Hájek projection of the gradient
All the point of working on $U_{N}^{*}$ and note directly on $U_{N}$ is that, conditionally on $X$, the $\bar{s}_{{\bf i}}$’s are independent because $\bar{s}_{\bf i}$ is $\sigma(X,Y_{\bf i})$-measurable and the edge weights $Y_{{\bf i}}$’s are independent conditionally on $X$. To prove a.n. for $U_{N}^{*}$ we apply Chatterjee’s result MR2294976 as in graham2017econometric.
@@
Now, as in MR2294976 and graham2017econometric, we consider independent Gaussian variables with the same first and second moments as the $R_{\bf i}$’s: let $(G_{\bf i})_{\bf i}$ be independent Gaussian variable with mean 0 and variances $(\sigma_{\bf i}^{2})_{\bf i}$. In particular, we observe that, conditionally on $X$, $n^{-1/2}\sum_{\bf i}G_{\bf i}$ is a ${\cal N}(0,1)$ random variable since $n^{-1}\sum_{\bf i}\sigma_{\bf i}^{2}=1$. As a consequence, in order to show that $n^{-1/2}\sum_{{\bf i}}R_{\bf i}\overset{d}{\to}{\cal N}(0,1)$ conditionally on $X$, it is enough to show that for all functions $f\in{\cal C}^{3}$ so that $\left\|f^{(r)}\right\|_{\infty}\leq L$ for all $r=0,1,2,3$, we have
@@
Let $f\in{\cal C}^{3}$ so that $\left\|f^{(r)}\right\|_{\infty}\leq L$ for all $r=0,1,2,3$. Our aim is to show that (B.13) holds. As in MR2294976 and graham2017econometric, we consider
@@ B.2.3 Uniform convergence of the Hessian over the compact set $B_{2}(\beta_{\star},1)$
In graham2017econometric; jochmans2018semiparametric, the authors apply Lemma 2.9 from newey1994large to prove the uniform convergence of the sequence of Hessian matrices of the loss function. However, the latter result requires the existence of a limit risk function that we don’t have. We therefore apply Corollary 2.2 from newey91. We recall this result adapted to our setup (we work with functions with values in $\mathbb{R}^{p\times p}$ unlike newey91).
Proposition 3 (Corollary 2.2 in newey91).
@@
The proof is a straightforward application of the one-dimensional result given in Corollary 2.2 in newey91 since one can write
@@ C.2 General results on orbits, $m_{\xi}$ and $M_{\xi}$ and the proof of Lemma 1
The proof of Lemma A.1 provided above is based on three ingredients: 1) convexity of $\hat{Q}_{n}$, 2) the uniform convergence in probability result over the compact set $\theta_{0}+2\varepsilon B_{2}$ derived from Lemma C.5 and, 3) the consistency result over the compact model $\theta_{0}+2\varepsilon B_{2}$ that follows from Lemma C.4. We start with the proof of the latter result which is an adaptation of Theorem 2.1 from newey1994large.
@@
Let $\Theta^{\prime}$ be a countable dense subset of $\Theta$. According to the diagonalization argument from Andersen1982 (recalled in Lemma C.8 below), there exists a sub-sequence $(\psi_{n})_{n}$ of $(\phi_{n})_{n}$ such that for all $\theta\in\Theta^{\prime},|\hat{Q}_{\psi_{n}}(\theta)-Q_{\psi_{n}}(\theta)|\overset{a.s.}{\to}0.$
@@
The proof of Lemma A.1 requires to revisit classical results from convex analysis and asymptotic statistics. We first start with Theorem 10.6 and Theorem 10.8 from rockafellar-1970a that need to be adapted to our setup.
@@
(Adapted from Theorem 10.6 in rockafellar-1970a) Let $\Theta$ be a non empty, open and convex set in $\mathbb{R}^{d}$. Let $(f_{n})_{n}$ and $(g_{n})_{n}$ be two sequences of convex functions defined on $\Theta$. We assume that there exists $\Theta^{\prime}\subset\Theta$ such that $\Theta\subset{\rm conv}(\bar{\Theta}^{\prime})$ and
@@
Since $(g_{n})_{n}$ is a sequence of convex functions satisfying (a) and (b) from Theorem 10.6 in rockafellar-1970a, it follows from the latter theorem that $(g_{n})_{n}$ is equi-Lipschitzian relative to $K$. Next, since $(f_{n}-g_{n})(\theta)\to 0$ for all $\theta\in\Theta^{\prime}$ and $(g_{n}(\theta))_{n}$ is uniformly (in $n$) bounded from above for all $\theta\in\Theta^{\prime}$ and there is some $\theta_{1}$ for which $(g_{n}(\theta_{1}))$ is bounded from below, then for all $\theta\in\Theta^{\prime}$, $(f_{n}(\theta))_{n}$ is also bounded from above and $(g_{n}(\theta_{1}))$ is bounded from below. Hence, it follows from Theorem 10.6 in rockafellar-1970a that $(f_{n})_{n}$ is equi-Lipschitzian relative to $K$. We conclude that the sequence of differences $(f_{n}-g_{n})_{n}$ is also equi-Lipschitzian relative to $K$. ∎
@@
(adapted from Theorem 10.8 from rockafellar-1970a) Let $\Theta$ be a non empty set in $\mathbb{R}^{d}$. Let $(f_{n})_{n}$ and $(g_{n})_{n}$ be two sequences of functions defined on $\Theta$. We assume that for all compact set $K$ in $\Theta$, there exists $L>0$ such that for all $n$ large enough and all $x,y\in K$,
@@
The next result is the diagonalization method used to prove Theorem II.1 in Andersen1982 that we reproduce here in our setup for the sake of completeness.
@@
(Theorem II.1 from Andersen1982)Let $\Theta$ be a non empty set. Let $(\hat{Q}_{n})_{n}$ be a sequence of random functions and $(Q_{n})_{n}$ be a sequence of (deterministic) functions, all defined on $\Theta$. We assume that for all $\theta\in\Theta$, $\hat{Q}_{n}(\theta)-Q_{n}(\theta)\overset{p}{\to}0$ as $n\to+\infty$. Then, for any countable subset $\Theta^{\prime}$ in $\Theta$ there exists an increasing sequence $(\psi_{n})_{n}$ of integers such that for all $\theta\in\Theta^{\prime}$,
@@
Proof of Lemma C.8: Denote for all $n$, $\hat{H}_{n}=\hat{Q}_{n}-Q_{n}$. Denote by $(x_{j})_{j\in{\mathbb{N}}}$ the sequence of all elements in $\Theta^{\prime}$. Since $(\hat{H}_{n}(x_{1}))_{n}$ tends in probability to 0, we can extract a sub-sequence $(\phi_{n,1})_{n}$ along which the converge is almost sure. Next, since $(\hat{H}_{\phi_{n,1}}(x_{2}))_{n}$ tends in probability to 0, it is possible to further construct a sub-sequence $(\phi_{n,2})_{n}$ along which the converge is almost sure. We repeat the argument and get that for all integers $k$, it is possible to construct a sequence $(\phi_{n,k})_{n}$ such that for all $i\in[k]$, $(\hat{H}_{\phi_{n,k}}(x_{i}))_{n}$ tends almost surely to 0. Once, this construction is done, we move to the diagonalization argument used in the proof of Theorem II.1 from Andersen1982: we construct a new sequence of integers, denoted by $(\psi_{n})_{n}$, by setting $\psi_{1}$ to be the first element in $(\phi_{n,1})_{n}$, then $\psi_{2}$ to be the second one in $(\phi_{n,2})_{n}$, etc.. Along this new sequence $(\psi_{n})_{n}$ we have for all $i\in{\mathbb{N}}$, $(\hat{H}_{\psi_{n}}(x_{i}))_{n}$ converges almost surely to 0. $\square$
@@ C.5 A result on the expectation of the empirical $L_{3}/L_{2}$ ratio and other probability results
Lemma C.12 (Example 11.9 from MR1652247).
@@ Appendix D Health data application
Results reported in Table D.1 show that this reform led physicians to increase their activity by between 9 and 14%, depending on the specification and the control group considered. wilner2025physician further show that the effect is more pronounced for young doctors, and they also find suggestive evidence that access to healthcare improved in the sense that doctors actually admitted new patients following the fee increase.

### v2

@@ B.2.1 Hájek projection of the gradient of the loss function at $\beta_{\star}$
The main difficulty in proving a CLT for $U_{N}$ is that it is not a sum of independent variables. However, we prove in this section, as in Graham (2017); Jochmans (2018), that it can be well approximated (in a sense given in Proposition 2 below) by its Hájek projection - which is a sum of independent variables.
@@
is achieved by $U_{N}^{*}$. Note that the $(\sigma(X,Y_{\bf i}))_{{\bf i}}$ are not independent in general, however, they are independent conditionally on $X$ because all edges weights are independent conditionally on $X$ in model assumption (1). That is the reason why we are considering the Hájeck projection conditionally to $X$. However, because of this conditioning, we cannot directly apply Lemma 11.10 from van der Vaart (1998) but we can adapt its proof to our context. We therefore follow the proof strategy of Lemma 11.10 from van der Vaart (1998) to show that $U_{N}^{*}$ is solution to (B.5).
@@ B.2.2 Asymptotic normality of the Hájek projection of the gradient
All the point of working on $U_{N}^{*}$ and note directly on $U_{N}$ is that, conditionally on $X$, the $\bar{s}_{{\bf i}}$’s are independent because $\bar{s}_{\bf i}$ is $\sigma(X,Y_{\bf i})$-measurable and the edge weights $Y_{{\bf i}}$’s are independent conditionally on $X$. To prove a.n. for $U_{N}^{*}$ we apply Chatterjee’s result Chatterjee (2006) as in Graham (2017).
@@
Now, as in Chatterjee (2006) and Graham (2017), we consider independent Gaussian variables with the same first and second moments as the $R_{\bf i}$’s: let $(G_{\bf i})_{\bf i}$ be independent Gaussian variable with mean 0 and variances $(\sigma_{\bf i}^{2})_{\bf i}$. In particular, we observe that, conditionally on $X$, $n^{-1/2}\sum_{\bf i}G_{\bf i}$ is a ${\cal N}(0,1)$ random variable since $n^{-1}\sum_{\bf i}\sigma_{\bf i}^{2}=1$. As a consequence, in order to show that $n^{-1/2}\sum_{{\bf i}}R_{\bf i}\overset{d}{\to}{\cal N}(0,1)$ conditionally on $X$, it is enough to show that for all functions $f\in{\cal C}^{3}$ so that $\left\|f^{(r)}\right\|_{\infty}\leq L$ for all $r=0,1,2,3$, we have
@@
Let $f\in{\cal C}^{3}$ so that $\left\|f^{(r)}\right\|_{\infty}\leq L$ for all $r=0,1,2,3$. Our aim is to show that (B.13) holds. As in Chatterjee (2006) and Graham (2017), we consider
@@ B.2.3 Uniform convergence of the Hessian over the compact set $B_{2}(\beta_{\star},1)$
In Graham (2017); Jochmans (2018), the authors apply Lemma 2.9 from Newey and McFadden (1994) to prove the uniform convergence of the sequence of Hessian matrices of the loss function. However, the latter result requires the existence of a limit risk function that we don’t have. We therefore apply Corollary 2.2 from Newey (1991). We recall this result adapted to our setup (we work with functions with values in $\mathbb{R}^{p\times p}$ unlike Newey (1991)).
Proposition 3 (Corollary 2.2 in Newey (1991)).
@@
The proof is a straightforward application of the one-dimensional result given in Corollary 2.2 in Newey (1991) since one can write
@@ C.2 General results on orbits, $m_{\xi}$ and $M_{\xi}$ and the proof of Lemma 1
The proof of Lemma A.1 provided above is based on three ingredients: 1) convexity of $\hat{Q}_{n}$, 2) the uniform convergence in probability result over the compact set $\theta_{0}+2\varepsilon B_{2}$ derived from Lemma C.5 and, 3) the consistency result over the compact model $\theta_{0}+2\varepsilon B_{2}$ that follows from Lemma C.4. We start with the proof of the latter result which is an adaptation of Theorem 2.1 from Newey and McFadden (1994).
@@
Let $\Theta^{\prime}$ be a countable dense subset of $\Theta$. According to the diagonalization argument from Andersen and Gill (1982) (recalled in Lemma C.8 below), there exists a sub-sequence $(\psi_{n})_{n}$ of $(\phi_{n})_{n}$ such that for all $\theta\in\Theta^{\prime},|\hat{Q}_{\psi_{n}}(\theta)-Q_{\psi_{n}}(\theta)|\overset{a.s.}{\to}0.$
@@
The proof of Lemma A.1 requires to revisit classical results from convex analysis and asymptotic statistics. We first start with Theorem 10.6 and Theorem 10.8 from Rockafellar (1970) that need to be adapted to our setup.
@@
(Adapted from Theorem 10.6 in Rockafellar (1970)) Let $\Theta$ be a non empty, open and convex set in $\mathbb{R}^{d}$. Let $(f_{n})_{n}$ and $(g_{n})_{n}$ be two sequences of convex functions defined on $\Theta$. We assume that there exists $\Theta^{\prime}\subset\Theta$ such that $\Theta\subset{\rm conv}(\bar{\Theta}^{\prime})$ and
@@
Since $(g_{n})_{n}$ is a sequence of convex functions satisfying (a) and (b) from Theorem 10.6 in Rockafellar (1970), it follows from the latter theorem that $(g_{n})_{n}$ is equi-Lipschitzian relative to $K$. Next, since $(f_{n}-g_{n})(\theta)\to 0$ for all $\theta\in\Theta^{\prime}$ and $(g_{n}(\theta))_{n}$ is uniformly (in $n$) bounded from above for all $\theta\in\Theta^{\prime}$ and there is some $\theta_{1}$ for which $(g_{n}(\theta_{1}))$ is bounded from below, then for all $\theta\in\Theta^{\prime}$, $(f_{n}(\theta))_{n}$ is also bounded from above and $(g_{n}(\theta_{1}))$ is bounded from below. Hence, it follows from Theorem 10.6 in Rockafellar (1970) that $(f_{n})_{n}$ is equi-Lipschitzian relative to $K$. We conclude that the sequence of differences $(f_{n}-g_{n})_{n}$ is also equi-Lipschitzian relative to $K$. ∎
@@
(adapted from Theorem 10.8 from Rockafellar (1970)) Let $\Theta$ be a non empty set in $\mathbb{R}^{d}$. Let $(f_{n})_{n}$ and $(g_{n})_{n}$ be two sequences of functions defined on $\Theta$. We assume that for all compact set $K$ in $\Theta$, there exists $L>0$ such that for all $n$ large enough and all $x,y\in K$,
@@
The next result is the diagonalization method used to prove Theorem II.1 in Andersen and Gill (1982) that we reproduce here in our setup for the sake of completeness.
@@
(Theorem II.1 from Andersen and Gill (1982))Let $\Theta$ be a non empty set. Let $(\hat{Q}_{n})_{n}$ be a sequence of random functions and $(Q_{n})_{n}$ be a sequence of (deterministic) functions, all defined on $\Theta$. We assume that for all $\theta\in\Theta$, $\hat{Q}_{n}(\theta)-Q_{n}(\theta)\overset{p}{\to}0$ as $n\to+\infty$. Then, for any countable subset $\Theta^{\prime}$ in $\Theta$ there exists an increasing sequence $(\psi_{n})_{n}$ of integers such that for all $\theta\in\Theta^{\prime}$,
@@
Proof of Lemma C.8: Denote for all $n$, $\hat{H}_{n}=\hat{Q}_{n}-Q_{n}$. Denote by $(x_{j})_{j\in{\mathbb{N}}}$ the sequence of all elements in $\Theta^{\prime}$. Since $(\hat{H}_{n}(x_{1}))_{n}$ tends in probability to 0, we can extract a sub-sequence $(\phi_{n,1})_{n}$ along which the converge is almost sure. Next, since $(\hat{H}_{\phi_{n,1}}(x_{2}))_{n}$ tends in probability to 0, it is possible to further construct a sub-sequence $(\phi_{n,2})_{n}$ along which the converge is almost sure. We repeat the argument and get that for all integers $k$, it is possible to construct a sequence $(\phi_{n,k})_{n}$ such that for all $i\in[k]$, $(\hat{H}_{\phi_{n,k}}(x_{i}))_{n}$ tends almost surely to 0. Once, this construction is done, we move to the diagonalization argument used in the proof of Theorem II.1 from Andersen and Gill (1982): we construct a new sequence of integers, denoted by $(\psi_{n})_{n}$, by setting $\psi_{1}$ to be the first element in $(\phi_{n,1})_{n}$, then $\psi_{2}$ to be the second one in $(\phi_{n,2})_{n}$, etc.. Along this new sequence $(\psi_{n})_{n}$ we have for all $i\in{\mathbb{N}}$, $(\hat{H}_{\psi_{n}}(x_{i}))_{n}$ converges almost surely to 0. $\square$
@@ C.5 A result on the expectation of the empirical $L_{3}/L_{2}$ ratio and other probability results
Lemma C.12 (Example 11.9 from van der Vaart (1998)).
@@ Appendix D Health data application
Results reported in Table D.1 show that this reform led physicians to increase their activity by between 9 and 14%, depending on the specification and the control group considered. Wilner and Choné (2025) further show that the effect is more pronounced for young doctors, and they also find suggestive evidence that access to healthcare improved in the sense that doctors actually admitted new patients following the fee increase.